# Memory in the Burst Occurrence of Repeating FRBs

### v1

@@ 4 Summary
In conclusion, we have analyzed the statistics of waiting times between bursts (above certain fluence threshold values $I$) of three repeating FRBs. Our observations reveal a notable memory effect in these events, wherein small waiting times tend to follow small ones, and large waiting times follow large ones. This clustering of FRBs has been quantified using four metrics: (i) the conditional distribution function, (ii) the conditional mean recurrence time, (iii) the conditional risk function, and (iv) the conditional residual time. Each of these metrics significantly deviates from their respective unconditional counterparts. The memory effect proves particularly beneficial as it enables us, as demonstrated in Fig. 5, to utilize relevant elapsed time data since last burst to achieve a more accurate burst estimation. For both repeating FRBs in their active periods and the apparent non-repeating FRBs that have been observed to burst only once, a unified description of waiting time is provided.
@@
All data analysed in this paper are included in references (Li et al., 2021b; Xu et al., 2022; Zhang et al., 2023; Zhou et al., 2022) separately. They are also openly available in Science Data Bank at https://doi.org/10.11922/sciencedb.01092 (Li et al., 2021b), https://cstr.cn/31253.11.sciencedb.08058 (Zhang et al., 2023), https://cstr.cn/31253.11.sciencedb.06762 (Zhou et al., 2022), and at the PSRPKU website, https://psr.pku.edu.cn/index.php/publications/frb20201124a/ (Xu et al., 2022).
Acknowledgements: We would like to thank Mingyu Ge and Sai Wang for useful discussions. The authors acknowledge the computing resources provided by National High Energy Physics Data Center and High Energy Physics Data Center, Chinese Academy of Science. S.-L.X. acknowledge the support by the National Natural Science Foundation of China (Grant No. 12273042), S.X. acknowledge the support by the National Natural Science Foundation of China (NO. 12303043), and C.C. acknowledge the support by the Science Foundation of Hebei Normal University (No. L2023B11), the National Natural Science Foundation of China (No. 12303045) and Natural Science Foundation of Hebei Province (No. A2023205020).

### v2

@@ 4 Summary
In conclusion, we have analyzed the statistics of waiting times between bursts (above certain fluence threshold values $I$) of three repeating FRBs. Our observations reveal a notable memory effect in these events, wherein small waiting times tend to follow small ones, and large waiting times follow large ones. This clustering of FRBs has been quantified using four metrics: (i) conditional distribution function, (ii) conditional mean recurrence time, (iii) conditional risk function, and (iv) conditional residual time. Each of these metrics generally deviates from their respective unconditional counterparts. The memory effect proves particularly beneficial as it enables us, as demonstrated in Fig. 5, to utilize relevant elapsed time data since the last burst to achieve a more accurate burst estimation. For both repeating FRBs in their active periods and the apparent non-repeating FRBs that have been observed to burst only once, a unified description of waiting time is provided. As a natural conclusion from the above results: Even if all FRBs are essentially repeating FRBs, the longer it has been since we have seen from those currently classified as one-off FRBs, the longer we may have to wait to see another burst, even if it is a repeating FRB. Conversely, the more recently observed one-off FRBs are more likely to be seen bursting again in a shorter period of time.
@@
All data analyzed in this article are included in references (Li et al., 2021b; Xu et al., 2022; Zhang et al., 2023; Zhou et al., 2022) separately. They are also openly available in the Science Data Bank at https://doi.org/10.11922/sciencedb.01092 (catalog ) 10.11922/sciencedb.01092 (Li et al., 2021b), https://cstr.cn/31253.11.sciencedb.08058 (catalog ) 10.57760/sciencedb.08058 (Zhang et al., 2023), https://cstr.cn/31253.11.sciencedb.06762 (catalog ) 10.57760/sciencedb.06762 (Zhou et al., 2022), and https://doi.org/10.6084/m9.figshare.19688854.v2 (catalog ) 10.6084/m9.figshare.19688854.v2 (Xu et al., 2022).
Acknowledgements: We thank Mingyu Ge, Sai Wang, and Xiaobo Li for useful discussions. S.-L.X. acknowledge the support of the National Natural Science Foundation of China (Grant No. 12273042), S.X. acknowledge the support by the National Natural Science Foundation of China (NO. 12303043) and C.C. acknowledge the support by the Science Foundation of Hebei Normal University (No. L2023B11), the National Natural Science Foundation of China (No. 12303045) and the Natural Science Foundation of Hebei Province (No. A2023205020). The authors acknowledge the computing resources provided by the National High Energy Physics Data Center and the High Energy Physics Data Center of the Chinese Academy of Science.